calculate the molarity and molality of 15% solution of H2So4 if Density is 1.020

Grade:12

2 Answers

huzaib khan
6 years ago
student-name Shikhar Chandra asked in ChemistryFind the molarity and molality of a 15% solution of H2SO4 ( density=1.020 g/cm3 and atomic mass = H=1, O=16, S=32amu) 22 Follow 0student-name Kanika answered this1349 helpful votes in Chemistry, Class XII-ScienceDear Student,Molarity can be calculated as:where %mass = 15 %d = 1.020 g/cm3MM = 98 g/molNow, Molality can be calculated with the help of Molarity, using the following formula:where, M= molarity = 1.56 Md = density = 1.020 g/cm3MM = Molar mass of solute = 98 g/mol
huzaib khan
6 years ago
Molarity can be calculated as:where %mass = 15 %d = 1.020 g/cm3MM = 98 g/molNow, Molality can be calculated with the help of Molarity, using the following formula:where, M= molarity = 1.56 Md = density = 1.020 g/cm3MM = Molar mass of solute = 98 g/mol